# Mini-Monkey: Alleviate the Sawtooth Effect by Multi-Scale Adaptive Cropping

### v1

@@ Abstract
Recently, there has been significant interest in enhancing the capability of multimodal large language models (MLLMs) to process high-resolution images. Most existing methods focus on adopting a cropping strategy to improve the ability of multimodal large language models to understand image details. However, this cropping operation inevitably causes the segmentation of objects and connected areas, which impairs the MLLM’s ability to recognize small or irregularly shaped objects or text. This issue is particularly evident in lightweight MLLMs. Addressing this issue, we propose Mini-Monkey, a lightweight MLLM that incorporates a plug-and-play method called multi-scale adaptive crop strategy (MSAC). Mini-Monkey adaptively generates multi-scale representations, allowing it to select non-segmented objects from various scales. To mitigate the computational overhead introduced by MSAC, we propose a Scale Compression Mechanism (SCM), which effectively compresses image tokens. Mini-Monkey achieves state-of-the-art performance among 2B-parameter MLLMs. It not only demonstrates leading performance on a variety of general multimodal understanding tasks but also shows consistent improvements in document understanding capabilities. On the OCRBench, Mini-Monkey achieves a score of 802, outperforming 8B-parameter state-of-the-art model InternVL2-8B. Besides, our model and training strategy are very efficient, which can be trained with only eight RTX 3090. The code is available at https://github.com/Yuliang-Liu/Monkey.
@@
1 Introduction
In recent years, the field of natural language processing (NLP) has demonstrated a significant paradigm shift, marked by a focus on the development of Large Language Models [77, 3, 64, 54] (LLMs). This shift has paved the way for the creation of multimodal large language models (MLLMs) capable of processing general vision-and-language understanding [32, 40, 2]. Researchers are actively exploring effective and efficient methods for integrating vision encoders with LLMs. Some methods, such as Flamingo [1], BLIP-2 [32], MiniGPT4 [79], and Qwen-VL [2] utilize a set of learnable queries to sample the image tokens and align the image tokens with Large Language Models. In contrast, other methods like LLaVA [41] and CogVLM [65] propose to use a linear layer to achieve this. Despite these achievements, detailed scene understanding was not achieved well by previous multimodal large language models due to the limited resolution to handle.
Figure 1: Sawtooth Effect caused by the cropping. (a) Input Image (b) Cropping strategy. (c) Overlapping Cropping Strategy. (d) Ours: Multi-scale adaptive crop strategy.
Recent efforts have attempted to tackle this issue by expanding the input resolution of the image. The cropping strategy is one of the most commonly used methods [39, 71, 35, 7, 58, 70]. There are many technical extensions to the simplest cropping strategy. For instance, Monkey [35] leverages the LoRA [22] into the vision encoder for learning detail-sensitive features from the sub-image. Although these methods have shown promising results, their performance still lags behind leading commercial models. To bridge this gap, InternVL 1.5 [7] employs a powerful vision encoder [8] to enhance visual representation and uses dynamic high resolution to scale up the resolution to 4K, significantly improving the performance.
@@ 1 Introduction
Despite the significant progress achieved by multimodal large language models, challenges in detailed scene understanding persist due to the cropping strategy. Cropping operations on images inevitably segment objects and connected areas, impairing the MLLM’s ability to recognize small or irregularly shaped objects, particularly in the context of document understanding. This strategy will introduce two types of semantic incoherence: 1) If an object or character is divided, it may not be recognized [24]. For instance, after cropping, the nose looks very much like a monkey, as shown in Fig. 1(b); 2) If a word or sentence is segmented, the semantic damage of the segmented word will be caused. For example, the word ‘Breakdown’ may be divided into ‘Break’ and ‘down’, causing semantic damage to the segmented word [44, 76]. For simplicity, we call this issue the sawtooth effect in this paper. A very straightforward idea is to adopt an overlapping cropping strategy to solve this issue, as presented in Fig. 1(c). However, as presented in our ablation studies Sec. 4.3, the overlapping cropping strategy introduces certain hallucinations that cause performance to decrease rather than increase. Moreover, this sawtooth effect is particularly evident in lightweight MLLMs, as discussed in Sec. 4.4. Larger MLLMs with enhanced comprehension capabilities can alleviate this issue to some extent.
In this paper, we propose Mini-Monkey, a lightweight multimodal large language model designed to mitigate the sawtooth effect caused by cropping strategies. Unlike existing methods [71, 7, 36] that directly crop input images, Mini-Monkey employs a plug-and-play method termed multi-scale adaptive crop strategy (MSAC), which enables effective complementation between features from different scales, as shown in Fig. 1(d). MSAC first performs a stratified operation on a pre-set group of grids according to the aspect ratios and the resolution of these grids. It then adaptively selects multiple aspect ratios from each stratified layer, ensuring that the same text is not split across different images. Multiple images will be generated based on the aspect ratios and processed by a pre-trained vision encoder to generate multi-scale visual representations. These representations are concatenated into a sequence and fused within the LLM to interact with each other. With the MSAC, Mini-Monkey adaptively generates multi-scale representations, allowing the model to select non-segmented object features from various scales. The MSAC may introduce some additional computational overhead. Therefore, we propose a Scale Compression Mechanism (SCM) for use in situations where there are restrictions on computational overhead. SCM is a training-free and parameter-free module to reduce the computational overhead. It utilizes the well-trained attention layers from the LLM to produce the attention weight and dropout token based on the attention weight.
Experiments have demonstrated the effectiveness of Mini-Monkey: 1) Mini-Monkey achieves state-of-the-art performance among 2B-parameter MLLMs in both general multimodal understanding and document understanding tasks. Especially, Mini-Monkey outperforms the state-of-the-art 2B-parameter method by an average of 1.7% across 13 benchmarks in terms of evaluation metrics; 2) Surprisingly, we find that Mini-Monkey achieves a score of 802 on the OCRBench, outperforming the 8B-parameter state-of-the-art model InternVL2-8B.
2 Related Works
@@ 2.1 Multimodal Large Language Models
In recent years, Large Language Models (LLMs) have made significant progress [77, 3, 64, 54, zeng2022glm,llama32024]. Drawing from this advancement, many efforts have been made to integrate a vision encoder into Large Language Models for vision-language understanding. A commonly employed approach is the Linear Projector method[40, 65, 79], which maps the output of the vision encoder to the same feature space as the text features of the Large Language Models. Some methods, such as QFormer [32], Perceiver Resampler [1, 2], or Abstractor [72], introduce a set of learnable queries to facilitate this integration. Despite notable progress, previous methods face challenges in handling the detailed scene understanding due to the limited resolution. To address this issue, recent works have primarily employed the following strategies: 1) Two vision encoders, one for processing high-resolution images and one for processing low-resolution images [67, 78, 20]. 2) Directly using visual encoders that support high-resolution input [37, 48]. 3) Using a cropping strategy to segment the high-resolution images into several low-resolution images [71, 36]. While these methods have effectively enhanced resolution, they still display shortcomings in document understanding, especially when compared to top commercial models. To close the gap, InternVL1.5 [7] utilizes a large vision encoder [8] and a dynamic high-resolution strategy to train on high-quality data. Concurrently, LLama3-V [63] employed a cropping strategy to enhance resolution, releasing several models with varying parameter counts, reaching up to 400 billion. Although LLama3-V and InternVL1.5 achieves promising results on several multimodal benchmarks, the cropping strategy used in it will inevitably result in semantic incoherence: 1) If an object or character is divided, it may not be recognized; 2) If the word or sentence is segmented, the semantic damage of the segmented word will be caused. For example, the word ‘Breakdown’ may be divided into ‘Break’ and ‘down’, causing semantic damage to the segmented word [44, 76]. This will limit it in the detailed scene understanding. Although some methods [44, 24] attempt to address this issue by introducing attention modules, they introduce additional parameters and require training this module from scratch. In contrast, our method is plug-and-play, requiring no additional parameters.
2.2 Visually-situated document understanding
The visually-situated document understanding is a task that comprehends rich text information in the images, including natural images [59, 60], documents images [53, 61, 51], charts images [50, 26], tables images [55, 6], etc. document understanding models can be broadly categorized into two types based on their reliance on OCR systems for text extraction: OCR-dependent methods and OCR-free methods. OCR-dependent methods use the text extracted from the OCR system to perform related document understanding tasks. For instance, LayoutLM v3 [25] learns the multimodal representations by unified text and image masking pre-training objectives. UDOP [62] develops a unified framework to learn and generate vision, text, and layout modalities together. On the contrary, OCR-free methods perform the document understanding tasks in an end-to-end manner without OCR input. Dount [28] directly generates textual elements based on the document images without OCR input. Pix2Struct [30] further leverages structure information by generating the HTML DOM tree for masked screenshots of web pages. Additionally, some methods incorporate large language models (LLMs) for enhanced document understanding. Ureader [71] presents a shape-adaptive cropping module to crop input images based on their aspect ratios. Similarly, TextMonkey [44] introduces enhancements for cross-window relations and a token resampler for document understanding. LayoutLLM [48] develops LayoutCoT for layout-aware supervised fine-tuning. StrucTexTv3 [49] introduces a lightweight multimodal large language model designed for perceiving and comprehending text-rich images.
@@ 2.3 Lightweight Multimodal Large Language Models
Due to the substantial computational costs associated with multimodal large language models (MLLMs), recent efforts have focused on developing more efficient models for rapid development and real-world applications. LLaVA-Phi [80] and Imp [57] leverage a lightweight language model combined with a vision encoder to create a lightweight large multimodal model. In this context, several researchers are exploring efficient architectural designs. For instance, MobileVLM [9] introduces a lightweight downsample projector to minimize resource usage during training and inference. Bunny [19] offers an efficient data compression technique to reduce the volume of pretraining data required. TinyGPT-V [75] employs a multi-stage training approach tailored for lightweight multimodal models. Similarly, MiniCPM presents a scalable training strategy aimed at producing an efficient lightweight multimodal large language model. Additionally, Vary-toy [68] supports high-resolution input, while InternVL 2 [7] enhances the performance of lightweight MLLMs through a dynamic high-resolution strategy. Despite these promising advancements, the state-of-the-art method faces limitations due to a sawtooth effect caused by the cropping strategy.
Figure 2: The overall architecture of Mini-Monkey. H-Attn represents high attention weight. L-Attn represents low attention weights. The tokens with low attention weights will be filtered. The shared LLM’s Layer represents using the block layer from LLM in SCM.
@@ 3 Mini-Monkey
The overall architecture is illustrated in Fig. 2. Mini-Monkey consists of an MSAC, a vision encoder, an MLP layer, a Scale Compression Mechanism, and a Large Language Model (LLM). Initially, Mini-Monkey generates multiple images through MSAC. These images are then processed by the vision encoder and MLP layer to extract image tokens. The Scale Compression Mechanism adjusts these image tokens based on the input question and forwards them to the LLM, which subsequently generates the final answers.
3.1 Multi-Scale Adaptive Crop Strategy
Previous state-of-the-art methods [36, 7] adopt the cropping-based strategy to expand the resolution of images and segment the high-resolution images into a set of sub-images. However, this cropping strategy will lead to a sawtooth effect. To address this issue, we introduce a multi-scale adaptive crop strategy (MSAC) that achieves the synergy between images with different scales to mitigate the semantic incoherence caused by the cropping strategy. As shown in Fig. 2 (b), we generate a pre-defined set of grids. The maximum of these grids is less than $Max_{num}$.
Then, we perform a stratified operation on these grids, which are divided into three sets according to their aspect ratios. We will select one aspect ratio for each layer. Different stratified layers provide different information to the model. The detailed layer $A_{d}$ is responsible for providing detailed information. It not only limits the maximum of the sub-image but also limits the minimum of the sub-image to make the image as large as possible to make the object in the image clearer. The images generated by this layer may have a semantic inconsistency. Therefore, we utilize the adaptive layer $A_{a}$ to synergize with the detailed layer, allowing the model to select non-segmented objects from various scales. The adaptive layer will adaptively generate an aspect ratio based on the detailed layer, ensuring that the cropping lines on the detailed layer and those on the adaptive layer do not overlap. This can be formulated as follows:
$C_{Id}\cap C_{Ia}=\emptyset\,.$ (1)
where $C_{Id}$ represents the cropping lines on the detailed layer. $C_{Ia}$ represents the cropping lines on the adaptive layer. Specifically, if the aspect ratio from the detailed layer is a multiple of that from the adaptive layer, we remove it from the adaptive layer and select a new ratio. This process ensures that the detailed and adaptive layers provide distinct semantic information and visual features for the model.
We also produce a global view of the image as a low resolution using an aspect ratio 1:1{1:1}, termed global layer. After obtaining the image from three layers, these images are sent to the vision encoder to extract the features and compress the visual tokens through the scale compression mechanism. Once the visual tokens are compressed, they will be fed into the large language model to conduct the multi-scale visual representations fusion and output the results.
Multi-Scale Visual Representations Fusion. Unlike the previous work [58] that simply concatenates multi-scale features along the dimension, our approach involves the fusion of multi-scale visual representations within a large language model (LLM). Within the LLM, these multi-scale visual representations interact with each other through self-attention. By fusing features from different scales, Mini-Monkey gains an enhanced ability to comprehend visual text information.
@@ 3.2 Scale Compression Mechanism
Although the proposed MSAC significantly enhances model performance, certain scenarios may impose computational requirements. To tackle this challenge, we introduce a parameter-free token compression method called the Scale Compression Mechanism (SCM), which is used to reduce the visual tokens. A well-trained LLM can effectively select the necessary visual features based on the input question. Consequently, SCM utilizes the first and second layers of the LLM to select visual tokens without generating any additional parameters. Specifically, as shown in Fig. 2 (c), the input visual token including $V_{d}\in\mathbb{R}^{L_{1}\times C}$, $V_{a}\in\mathbb{R}^{L_{2}\times C}$, and $V_{g}\in\mathbb{R}^{L_{3}\times C}$, and the textual token $T_{t}\in\mathbb{R}^{T\times C}$ will be sent into an LLM’s Layer. $V_{d}$ represents the tokens from detailed layer. $V_{a}$ represents the tokens from adaptive layer. $V_{g}$ represents the tokens from global layer. We choose the visual token from the adaptive layer, global, and textual token to attend to the visual token from the detailed layer. The LLM’s Layer will output an attention map. The calculation of the attention can be formulated as follows:
$\mathbf{Attn_{w}}={\tt softmax}(\frac{(Cat(V_{a},V_{g},T_{t})+{\rm PE}(Cat(V_{a},V_{g},T_{t})))(V_{t}+{\rm PE}(V_{t}))^{T}}{\sqrt{D}})\,.$ (2)
Where $\rm PE$ represents the position encoding and $D$ denotes the dimension of the LLM, we proceed to obtain the attention map $\mathbf{Attn_{w}}\in\mathbb{R}^{T\times L}$. $Cat()$ represents the sequence concatenation operation. Due to the lower information density of tokens from detailed layers, we primarily focus on compressing these tokens. In contrast, visual tokens from adaptive and global layers provide the LLM with complete spatial information. After obtaining the attention map, we select the top $K$ visual features from detailed layers based on this map. Compared to FastV [5], SCM is more targeted by using tokens with high relative information density to compress tokens with low information density. The ablation study in Sec. 4.3 demonstrates the effectiveness of SCM.

### v3

# Mini-Monkey: Alleviating the Semantic Sawtooth Effect for Lightweight MLLMs via Complementary Image Pyramid

@@ Abstract
[…]
@@
Recently, Large Language Models (LLMs) [73, 4, 58, 53] have received significant attention for their robust text understanding and generation capabilities. Researchers are actively exploring ways to integrate vision encoders into LLMs to upgrade them to multimodal large language models (MLLMs) [33, 41, 2]. Some approaches employ a Q-former [1, 33], while others [42, 60] use linear projection. Despite the promising results, they are constrained to processing low-res images, which limits their ability to execute detailed scene analysis.
Figure 1: Comparison of different image cropping strategies. (a) Input image; (b) Non-overlapping cropping; (c) Overlapping cropping; (d) Ours: complementary image pyramid.
To address this limitation, much recent effect aims to enable MLLMs to process high-res images. One straightforward solution is to adopt a visual encoder that can tackle high-res images. However, developing a high-quality visual encoder demands substantial training resources [2, 7]. An alternative, more resource-efficient strategy is the non-overlapping cropping [37, 40, 67, 36, 8], which splits a high-res image into a set of low-res sub-images.
@@
While the non-overlapping cropping strategy has shown promising results, it inevitably cuts off objects and connected regions, rendering difficulty for the MLLM in recognizing small or irregularly shaped objects due to semantic discontinuity, particularly in the context of document understanding. This mainly leads to two types of consequences: i) semantic ambiguity: if an object or character is divided, it may be misidentified; the nose of guinea pig in Fig. 1(b) looks much like a monkey after cropping, for instance; 2) semantic damage: if a word or sentence is segmented, the meanings of the segmented word will be changed completely; if the word ‘breakdown’ is divided into ‘break’ and ‘down’, the segmented words will have nothing to do with the original one [45, 72]. For simplicity, we call these phenomena the semantic sawtooth effect in this paper. To alleviate this effect, a rather straightforward idea is to adopt overlapping cropping. However, this strategy will result in the processing of much duplicate information, as presented in Fig. 1(c). This redundancy could potentially cause hallucinations in MLLMs. Moreover, it can even deteriorate performance according to our ablation studies in Sec. 3.3. Additionally, the semantic sawtooth effect can be observed more evidently in lightweight MLLMs. Larger MLLMs with enhanced comprehension capabilities and feature extraction capabilities often can alleviate this issue to some extent. Even when the object is segmented, these models can understand the objects through their powerful feature extraction.
To alleviate the semantic sawtooth effect more explicitly, we propose a plug-and-play approach, termed Complementary Image Pyramid (CIP). CIP can be easily integrated into a variety of cropping-based MLLMs, allowing them to tackle high-res images with reduced semantic sawtooth effect. CIP dynamically constructs an image pyramid that provides complementary semantic features for the MLLMs, enabling it rich acquire semantics at all levels. If object semantics are lost at one scale, they can be compensated by those from another scale. Different from previous work [45, 25] that addresses this issue by modifying the architecture of the model, our approach focuses on enriching the image semantics per se. Consequently, CIP can be easily integrated into a variety of MLLMs, allowing them to tackle high-res images with reduced semantic sawtooth effect. Considering that the CIP introduces some additional computational overheads, we further propose a Scale Compression Mechanism (SCM) for use in situations with limited computational resources. The SCM is both training-free and parameter-free. It leverages the well-trained attention layers of the LLM and the multi-scale information to generate attention weights, which in turn are used to compress redundant tokens. Utilizing the proposed CIP and SCM, we introduce a lightweight MLLM, Mini-Monkey.
Our experiments demonstrate the effectiveness of the proposed method: 1) 2B-parameter Mini-Monkey outperforms the InternVL2-2B by an average of $2.4\%$ across 17 benchmarks in terms of evaluation metrics; 2) Mini-Monkey achieves a score of 806 on the OCRBench, outperforming the 8B-parameter model InternVL2-8B by $12$ score. Moreover, we observe that directly fine-tuning well-performing pre-trained MLLM does not enhance, but rather degrades its performance. In contrast, fine-tuning with CIP can facilitate the training process to improve performance. In conclusion, the contributions of this work can be summarized as follows:
  • CIP: a plug-and-play complementary image pyramid designed to alleviate the semantic sawtooth effect for multimodal large language models;
  • Mini-Monkey: a lightweight, effective, and training-efficient multimodal large language model that integrates the complementary image pyramid and the scale compression mechanism;
  • Our method achieves promising results on 8 general multimodal understanding benchmarks and 9 document understanding benchmarks, demonstrating the benefits of alleviating the semantic sawtooth effect.
1 Related Work
@@ 1.1 Multimodal Large Language Models
In recent years, Large Language Models (LLMs) have made significant progress [73, 4, 58, 53]. Drawing from this advancement, many efforts have been made to integrate a vision encoder into Large Language Models for vision-language understanding. A commonly employed approach is the linear projector method[42, 60], which maps the output of the vision encoder to the same feature space as the text features of the Large Language Models. Some methods, such as Q-Former [33], Perceiver Resampler [1], or Abstractor [68], introduce a set of learnable queries to facilitate this integration. Despite these notable advances, previous methods often struggle with detailed scene understanding due to limitations in resolution.
To address this issue, recent research has adopted two primary strategies: 1) The direct use of visual encoders that support high-res input. Some methods [62, 22] utilize two vision encoders, one for processing high-res images and another for low-res images. Others [59, 46] leverage a single vision encoder. However, these methods require additional training data and parameters, or processing attention over high-res images significantly increases computational demands. 2) An alternative, more resource-efficient method is the cropping strategy, which divides the high-res image into multiple lower-resolution sub-images for processing. Some methods apply a fixed-size cropping scheme [36, 37]. Others adopt a dynamic cropping approach [67, 8, 15] to keep the aspect ratio of the original image. Although the cropping strategy achieves promising results on several multimodal benchmarks, it will inevitably result in a semantic sawtooth effect: 1) If an object or character is divided, it may not be recognized; 2) If the word or sentence is segmented, the semantic damage of the segmented word will be caused. For example, the word ‘Breakdown’ may be divided into ‘Break’ and ‘down’, causing semantic damage to the segmented word. This will limit the model’s ability to understand the detailed scene. Although some methods [45, 25] attempt to address this issue by introducing new modules, they introduce additional parameters to the original model and require training this module from scratch. In contrast, the proposed CIP is designed to be seamlessly integrated without introducing additional parameters, offering a plug-and-play solution. A recent work [55] also attempts to utilize the multi-scale image to enhance the model’s capacity. However, they employ a fixed resolution, lacking the ability to dynamically generate complementary semantics based on the resolution of input images. On the contrary, The CIP dynamically constructs an image pyramid that provides complementary semantic features after cropping for the MLLMs, enabling them to richly acquire semantics at all levels.
@@ 1.2 Lightweight Multimodal Large Language Models
Due to the substantial computational costs associated with multimodal large language models (MLLMs), some recent efforts have focused on developing more efficient models for rapid development and real-world applications. For instance, LLaVA-Phi [75] and Imp [54] integrate a lightweight large language model with a vision encoder to develop a powerful multimodal system. MobileVLM [10] further conserves resources by integrating a lightweight downsampling projector that reduces the number of visual tokens. Bunny [21] advances efficiency through an effective data compression technique, which minimizes the required pretraining dataset. TinyGPT-V [71] adopts a multi-stage training process specifically designed for lightweight multimodal models. The above models support only low-res input. To improve detailed scene understanding in lightweight MLLMs, one of the most commonly used methods is the cropping strategy. For instance, MiniCPM-V series [66] d employs an adaptive visual encoding method to accommodate high-res images of varying aspect ratios. Similarly, InternVL2-2B [8] enhances the performance of lightweight MLLMs by adopting a dynamic high-res cropping strategy. Despite these advancements, the cropping strategy will introduce a semantic sawtooth effect, which significantly limits the performance of lightweight multimodal large language models. Larger MLLMs with enhanced comprehension capabilities can alleviate this issue to some extent, as discussed in Sec. 3.4.
Figure 2: Overall architecture of Mini-Monkey. H-Attn represents high attention weights. L-Attn represents low attention weights. The tokens with low attention weights will be filtered. The shared LLM’s Layer represents using the block layer from LLM in SCM.
@@ 2 Mini-Monkey
The overall architecture is illustrated in Fig. 2. Mini-Monkey consists of a CIP, a vision encoder, an MLP layer, a Scale Compression Mechanism, and a Large Language Model (LLM). Initially, CIP dynamically generates an image pyramid based on the resolution of input images. Then, we divide these images into a set of sub-images. These sub-images are then processed by the vision encoder and MLP layer to extract image tokens. The Scale Compression Mechanism adjusts these image tokens and forwards them to the LLM, which subsequently generates the final answers.
2.1 Complementary Image Pyramid
Existing cropping strategy [36, 8] directly divides the high-res images into a set of sub-images that will lead to a semantic sawtooth effect. To address this issue, we propose a plug-and-play method, termed complementary image pyramid (CIP), to promote synergy among images at varying scales to alleviate the semantic sawtooth effect. The process of CIP is shown in Fig. 2 (b).
Grouping Operation. We first generate a set of pre-defined aspect ratios. During testing, we limit the maximum number of tiles to 24. A greater number of tiles corresponds to images with higher resolution. These aspect ratios are then categorized into three groups through a grouping operation, including a detailed group $D_{g}$, an adaptive group $A_{g}$, and a global group $G_{g}$. The classification is based on the following criteria: (1) Aspect ratios that result in more than 9 tiles being allocated to the detailed group, enabling the largest possible image size and thus a clearer depiction of the objects within. (2) For aspect ratios producing between 3 and 8 tiles, we classify them into the adaptive group, which is responsible for enhancing the fine details at the borders of the crops. (3) The 1:1 aspect ratio is designated to the global group, providing a low-res, comprehensive view of the whole image. The grouping operation generates three groups of different aspect ratios for generating the image.
Dynamically Generating Images. After the grouping operation, we will select one aspect ratio from each group to produce a total of three images. For the detailed group, we calculate the aspect ratio of the input image and then compare it with the aspect ratios within the detailed group by calculating the absolute differences. The aspect ratio that best matches the input image, denoted as $D_{h},D_{w}$, where $D_{h}$ is the number of height tiles and $D_{w}$ is the number of width tiles, is then used to resize the image and divide the resized image into a series of sub-images. After obtaining the detailed image, the adaptive group will dynamically generate an aspect ratio based on $D_{h},D_{w}$, ensuring that the cropping lines on the detailed group and those on the adaptive group do not overlap. To achieve this, we ensure that the aspect ratios of the adaptive and detailed groups are not integer multiples of one another. Consequently, we will dynamically remove any aspect ratios that are exact multiples of $D_{h},D_{w}$. This can be formulated as follows:
$\forall k\in\mathbb{Z},\,D_{h}\neq k\cdot A_{h}\quad\text{and}\quad D_{w}\neq k\cdot A_{w}$ (1)
where $A_{h}$ and $A_{w}$ denote the height and width components of the aspect ratios in the adaptive group, respectively. Then, we will resize the image by selecting the ratio closest to the aspect ratio of the original image from the remaining aspect ratio.
Due to the independent processing of sub-images in the vision encoder, previous cropping-based MLLMs have not facilitated feature interaction across different sub-images. In our method, adaptive group employ a distinct aspect ratio to partitioning windows compared to detailed group, thereby simulating cross-window feature interactions within the detailed group. The detailed image provides detailed information about the image. The adaptive image enhances the fine features at the cropping boundaries of the detailed image. The global image captures the global information of the entire image for the model. Different from previous method [45, 25], the proposed CIP alleviates the semantic sawtooth effect from the perspective of the image, bringing several advantages: (i) it is plug-and-play, requiring no additional parameters; (ii) it seamlessly integrates with existing MLLMs that utilize cropping strategies, leading to consistent performance improvements; and (iii) it can be utilized without training and its effectiveness can be further improved through fine-tuning.
@@ 2.2 Scale Compression Mechanism
Although the proposed CIP significantly enhances model performance, certain scenarios may restrict the level of computational resources available. To tackle this challenge, we introduce a parameter-free token compression method called the Scale Compression Mechanism (SCM), which is used to reduce the visual tokens, as shown in Fig. 2 (c). The detailed group provides tokens with lower information density, whereas the adaptive and global groups yield tokens that are more information-dense. Therefore, we primarily focus on compressing the tokens from the detailed group. Specifically, a well-trained LLM from MLLM can effectively select the necessary visual features based on the input question. Consequently, SCM utilizes the first and second layers of the LLM to select visual tokens without generating any additional parameters. The input visual token including $\mathbf{V_{d}}\in\mathbb{R}^{L_{1}\times C}$, $\mathbf{V_{a}}\in\mathbb{R}^{L_{2}\times C}$, and $\mathbf{V_{g}}\in\mathbb{R}^{L_{3}\times C}$, and the textual token $\mathbf{T_{t}}\in\mathbb{R}^{T\times C}$ will be sent into an LLM’s Layer. $\mathbf{V_{d}}$ represents the tokens from the detailed group. $\mathbf{V_{a}}$ represents the tokens from the adaptive group. $\mathbf{V_{g}}$ represents the tokens from the global group. Notable, we reuse the layer of the LLM as this LLM’s Layer. The LLM’s Layer will output an attention map. We use the visual token from the adaptive group, global group, and textual token to attend to the visual token from the detailed group. The calculation of the attention can be formulated as follows:
$\mathbf{Q}={\tt cat}(\mathbf{V_{a}},\mathbf{V_{g}},\mathbf{T_{t}})\,$, (2)
$\mathbf{Attn_{w}}={\tt softmax}(\frac{\mathbf{Q}+{\rm PE}(\mathbf{Q}))(\mathbf{V_{d}}+{\rm PE}(\mathbf{V_{d}}))^{T}}{\sqrt{D}})\,.$ (3)
where $\rm PE$ represents the position encoding and $D$ denotes the dimension of the LLM. $Cat()$ represents the sequence concatenation operation. After computing the attention mechanism, we average the first dimension of the attention map $\mathbf{Attn_{w}}\in\mathbb{R}^{(L_{2}+L_{3}+T)\times L_{1}}$ to obtain a weight vector $\mathbf{W_{a}}\in\mathbb{R}^{L_{1}}$. Subsequently, we select the top $K$ visual features from detailed layers based on this weight vector $\mathbf{W_{a}}$. These selected tokens, along with tokens from the adaptive group, global group, and textual token, are input into the LLM to generate the results. Compared to FastV [6], SCM works in conjunction with the CIP and is more targeted by using tokens with high relative information density to compress tokens with low information density.